OpenStudy (anonymous):

how to calculate the value of X, Y and Z in a sphere where by Y is perpendicular to XZ(i mean Y start from north pole to source pole. is it right to calculate the value of r as r=sqrt(x^2 + y^2) or r=sqrt(p^2-Z^2) will this be equal to r=sqrt(x^2 + z^2

OpenStudy (amistre64):

you might want to look into spherical coordinants

OpenStudy (anonymous):

will Z=Rcos(theta)sin(phi)? how can i find Z=Rxyproj*sin(phi) or Z=Rxyproj*sin(theta) Rxyproj is the projected radius of sphere to xyplane and R the radius of sphere. Y is perpendicular to ZX plane and pointing north pole to south pole

OpenStudy (amistre64):



OpenStudy (amistre64):

\[\rho=radius\]\[r=pcos\,\gamma\]\[x=rcos\,\alpha\]\[z=r\,cos\beta\]

OpenStudy (amistre64):

we can reduce this to 2 angles by stating\[z=rsin\,\alpha\]instead

OpenStudy (amistre64):

i see know what to correlate alpha with gamma tho. and im not sure if i really know what it is youre asking either

OpenStudy (amistre64):

oh, and \(\large y=\rho sin\,\gamma\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

for your explanation and solution Godbless you

OpenStudy (anonymous):

is it right to calculate the value of r as r=sqrt(x^2 + y^2) or r=sqrt(p^2-Z^2) will this be equal to r=sqrt(x^2 + z^2

OpenStudy (amistre64):